Find the value of y, given the system: (X+ Y+ Z= 6 x- y+ Z= 2 2X+ y-Z=1 O a) y=0 Ob) y=1 Oc) y = 2 Od) y=3
Answer #1

We have U +2 = 6 1 + 2 2 2x + y - Z=1 - from equation De we have x + y + z = 6 x + z fyz 6 K+2 = 6-7 - © from equation ® , wei have tried my best to solve this problem but still if you have any doubt regarding this question then please comment